# Charge radii and their deformation correlation for even-$Z$ nuclei in deformed relativistic Hartree-Bogoliubov theory in continuum

### v1

@@ Abstract
The systematics are investigated for the charge radii of the even-$Z$ nuclei with $8\leqslant Z\leqslant 120$ calculated by the deformed relativistic Hartree-Bogoliubov theory in continuum (DRHBc) with the functional PC-PK1, and their deformation correlation is explored. The available data of the charge radius are reproduced with a root-mean-square deviation $\sigma=0.033$ fm. In particular, for the nuclei between the closed shells, the descriptions of the charge radii are remarkably improved by including the deformation. Taking molybdenum isotopes as examples, both the evolutions of the charge radius and deformation are well reproduced. It is found that while the charge radius typically increases with the deformation, there also exist different cases. For example, in 346Sg, the charge radius of the deformed ground state is smaller than the one of the spherical state, and the largest binding energy does not necessarily correspond to the smallest charge radius.
@@ I Introduction
The relativistic density functional theory (RDFT) has been proved to be a powerful tool in nuclear physics by its successful applications in describing many nuclear phenomena Meng (2016). Due to its compelling advantages, such as the automatic inclusion of the spin-orbit coupling Koepf and Ring (1991); Ren and Zhao (2020) and the reasonable description of its isospin dependence Sharma et al. (1995), the new saturation mechanism by the competition between the scalar and vector densities Walecka (1974), the natural explanation of the pseudospin symmetry in the nucleon spectrum Ginocchio (1997); Meng et al. (1998a); Liang et al. (2015) and the spin symmetry in antinucleon spectrum Liang et al. (2015); Zhou et al. (2003a); He et al. (2006), and the self-consistent treatment of the nuclear magnetism Koepf and Ring (1989, 1990), the RDFT has continuously attracted worldwide attentions in recent decades Ring (1996); Vretenar et al. (2005); Meng et al. (2006); Nikšić et al. (2011); Meng et al. (2013); Meng and Zhou (2015); Zhou (2016); Meng (2016); Shen et al. (2019); Meng and Zhao (2021). Based on the RDFT, efforts have been made for the theoretical descriptions of the charge radii and the improvement of the accuracy An et al. (2020); Perera et al. (2021); An et al. (2022, 2024a).
@@
Except for doubly-magic nuclei which exhibit spherical shape, most atomic nuclei are deformed. Based on the RDFT and treating the effects of deformation, pairing correlations and continuum simultaneously, the deformed relativistic Hartree-Bogoliubov theory in continuum (DRHBc) was developed Zhou et al. (2010); Li et al. (2012a). Inheriting the advantages of the RCHB theory and including the deformation degrees of freedom, the DRHBc theory has been successfully applied in investigating the halo structures in B Yang et al. (2021); Sun (2021), C Sun et al. (2018, 2020); Wang et al. (2024), Ne Zhong et al. (2022); Pan et al. (2024), Na Zhang et al. (2023a), Mg Zhang et al. (2023b); An et al. (2024b), and Al Zhang et al. (2024a) isotopes, dripline locations In et al. (2021); Zhang et al. (2021); Pan et al. (2021); He et al. (2021, 2024), nuclear shapes and sizes Pan et al. (2019); Choi et al. (2022); Kim et al. (2022); Guo et al. (2023); Mun et al. (2023), nuclear stability and radioactive decays Xiao et al. (2023); Choi et al. (2024); Zhang et al. (2024b), as well as other applications and extensions Zhang et al. (2019); Sun and Zhou (2021a, b); Sun et al. (2022); Sun and Meng (2022); Zhang et al. (2022b, 2023c); Zhao et al. (2023); Mun et al. (2024); Zhang et al. (2024c); Pan et al. (2025). Based on the DRHBc theory, a nuclear mass table including both deformation and continuum effects is under construction Zhang et al. (2020); Pan et al. (2022). Up to now, the even-$Z$ part of the DRHBc mass table has been completed, and the ground-state properties, including the binding energy, charge radius and quadrupole deformation, have been published Zhang et al. (2022a); Guo et al. (2024a).
In this work, the systematics of the charge radii for even-$Z$ nuclei based on the DRHBc theory with PC-PK1 are investigated. In Sec. II, a brief framework of the DRHBc theory is introduced. The numerical details in the DRHBc mass table calculations are provided in Sec. III. The systematics of the charge radii and their deformation correlation are discussed in Sec. IV. A summary is given in Sec. V.
@@ II Theoretical framework
where $\braket{r^{2}}_{p}$ is the root-mean-square (rms) radius of proton.
@@ IV.1 Systematics of the charge radii
Figure 1(b) shows the charge radius differences between the DRHBc Guo et al. (2024a) and RCHB Xia et al. (2018) calculations, $R_{\text{ch}}^{\text{(DRHBc)}}-R_{\text{ch}}^{\text{(RCHB)}}$, for even-$Z$ nuclei. The rms difference is $\sigma=0.055$ fm. For nuclei near the regions with either proton or neutron magic numbers, the differences are negligible. Remarkable differences appear in the middle of the shell, for example, the regions with $N$ between 50 and 82, 82 and 126, 126 and 184, as well as 184 and 258. Among 4693 nuclei, there are 2859 nuclei with differences larger than 0.01 fm, i.e., 2783 nuclei with $R_{\text{ch}}^{\text{(DRHBc)}}>R_{\text{ch}}^{\text{(RCHB)}}$ and only 76 nuclei with $R_{\text{ch}}^{\text{(DRHBc)}}<R_{\text{ch}}^{\text{(RCHB)}}$. This indicates that the deformation typically leads to a larger radius.
@@ IV.2 Evolution of the charge radius with the neutron number
To investigate the evolution of the charge radius with respect to the neutron number and the roles played by the deformation in detail, the Mo ($Z=42$) isotopes, which exhibit rich information on shape evolution as shown in Fig. 1(c), are taken as examples. The charge radii of Mo isotopes calculated by the DRHBc theory from the proton drip line at $N=40$ to the neutron drip line at $N=112$ are shown in Fig. 2(a) with red diamonds. Since the isotopes with $54\leqslant N\leqslant 68$ exhibit shape coexistence, their prolate and oblate minima are also presented with green and orange triangles, respectively. For comparison, the RCHB results Xia et al. (2018) are shown with blue circles, the data available Angeli and Marinova (2013); Li et al. (2021) are shown with black squares, and the empirical formula $r_{0}A^{1/3}$ is shown with the dashed line, where $r_{0}=0.956$ fm is determined by the data at $N=50$. In Fig. 2(a), the empirical formula roughly reproduces the general trend of the data. For the isotopes near $N=50$ with spherical or near-spherical shapes, both the RCHB and DRHBc calculations reproduce the data well, including the kink at $N=50$ corresponding to the shell closure. For the isotopes with $N>52$, the data are underestimated by RCHB but well reproduced by DRHBc. For the isotopes with shape coexistence, $58\leqslant N\leqslant 66$, the charge radii corresponding to the prolate minima in the DRHBc calculations better reproduce the data. The rms deviation from the data is $\sigma=0.030$ fm for DRHBc, in comparison with $\sigma=0.062$ fm for RCHB. In short, taking into account the deformation effect remarkably improves the description of charge radii for Mo isotopes.
@@ IV.3 Deformation dependence of the charge radius
Figure 4: (Color online) Total energies $E_{\text{tot}}$ [(a), (b) and (c)] and charge radii $R_{\text{ch}}$ [(d), (e) and (f)] as functions of quadrupole deformation $\beta_{2}$ of the Mo isotopes with $N=52,58,68$ from the constrained DRHBc calculations with PC-PK1. The ground-state deformations are denoted by squares. The empirical formula for charge radius, $(1+5\beta^{2}/4\pi)R_{\text{sph}}$, is shown by the gray dashed line.
To further investigate the deformation dependence of the charge radius, the shape coexistence and the competition between the prolate and oblate minima, taking the Mo isotopes with $N=52,58$ and 68 as examples, the potential energy curves (PECs), as well as the evolutions of the charge radii with the quadrupole deformation are shown in Fig. 4. By assuming a uniform nuclear density and neglecting the higher orders of deformation, the empirical formula for the charge radius as a function of the quadrupole deformation parameter $\beta_{2}$ Bohr and Mottelson (1998),
$R_{\text{ch}}(\beta_{2})=\left(1+\frac{5}{4\pi}\beta_{2}^{2}\right)R_{\text{ch}}(0),$ (4)
is also shown in Fig. 4.
@@
For the isotope with $N=68$, the ground state corresponds to the oblate minimum with $\beta_{2}=-0.23$, and there is a spherical minimum with the excitation energy 0.70 MeV, as shown in Fig. 4(c). For the neighboring isotopes with $N=69$ and 70, the ground states will move to the near-spherical and spherical shapes, respectively. The charge radius increases monotonically with the absolute value of the quadrupole deformation parameter, as shown in Fig. 4(f). It seems that the charge radius always increases with the deformation.
@@
It is interesting to check whether the charge radii may decrease with the deformation $\beta_{2}$. In Fig. 5, the PEC, as well as the evolution of the charge radius with the quadrupole deformation are shown for 346Sg. In Fig. 5(a), there are two minima of the energy with $\beta_{2}=-0.15$ and $0.17$, and the ground state corresponds to the prolate one. The spherical result is a saddle point on the PEC, and its energy is 6.00 MeV higher than the ground state. In Fig.  5(b), there are two minima of the charge radius with $\beta_{2}=-0.15$ and $0.10$, which do not correspond to the minima in Fig. 5(a). Between these two minima of the charge radius, the charge radius increases with the decrease of the absolute value of the quadrupole deformation parameter. These results demonstrate that one cannot simply conclude the charge radius as a monotonically increasing function of the quadrupole deformation parameter, and the largest binding energy does not necessarily correspond to the lowest charge radius.
@@ V Summary
In the present, only the results of even-$Z$ nuclei are available Guo et al. (2024a). A complete DRHBc mass table including odd-$Z$ nuclei will appear soon, enabling a more comprehensive investigation on the systematics of the charge radii and their deformation correlation. More experimental data, including charge radii and deformations, will deepen our understanding on nuclear structures, especially the data distinguishing prolate from oblate shapes, which will provide valuable constraints in examining existing nuclear models. As the DRHBc theory is based on the axial symmetry approximation, the effects of the triaxial deformation Zhang et al. (2023d) and the deformations with higher orders on the charge radius are issues worthy of further investigation.

### v2

@@ Abstract
The systematics are investigated for the charge radii of the even-$Z$ nuclei with $8\leqslant Z\leqslant 120$ calculated by the deformed relativistic Hartree-Bogoliubov theory in continuum (DRHBc) with the functional PC-PK1, and their deformation correlation is explored. The available data of the charge radius are reproduced with a root-mean-square deviation $\sigma=0.033$ fm. In particular, for the nuclei between the closed shells, the descriptions of the charge radii are remarkably improved by including the deformation. Taking molybdenum isotopes as examples, both the evolutions of the charge radius and deformation are well reproduced. It is found that while the charge radius typically increases with the deformation, there also exist different cases. For example, in 346Sg, the charge radius of the deformed ground state is smaller than the one of the spherical state, and the largest binding energy does not necessarily correspond to the smallest charge radius. The increase or decrease of the charge radii with deformation is related to specific shell structures, particularly the key single-particle levels near the Fermi energy.
@@ I Introduction
The relativistic density functional theory (RDFT) has been proved to be a powerful tool in nuclear physics by its successful applications in describing many nuclear phenomena Meng (2016). Due to its compelling advantages, such as the automatic inclusion of the spin-orbit coupling Koepf and Ring (1991); Ren and Zhao (2020) and the reasonable description of its isospin dependence Sharma et al. (1995), the new saturation mechanism by the competition between the scalar and vector densities Walecka (1974), the natural explanation of the pseudospin symmetry in the nucleon spectrum Ginocchio (1997); Meng et al. (1998a); Liang et al. (2015) and the spin symmetry in antinucleon spectrum Liang et al. (2015); Zhou et al. (2003a); He et al. (2006), and the self-consistent treatment of the nuclear magnetism Koepf and Ring (1989, 1990), the RDFT has continuously attracted worldwide attentions in recent decades Ring (1996); Vretenar et al. (2005); Meng et al. (2006); Nikšić et al. (2011); Meng et al. (2013); Meng and Zhou (2015); Zhou (2016); Meng (2016); Shen et al. (2019); Meng and Zhao (2021). Based on the RDFT, efforts have been made for the theoretical descriptions of the charge radii and the improvement of the accuracy An et al. (2020); Perera et al. (2021); An et al. (2022, 2024a). The descriptions of the charge radius by the macroscopic-microscopic model and the nonrelativistic density functional theory can be found in Refs. Wang and Li (2013); Gaidarov et al. (2014); Bassem and Oulne (2015); Sarriguren (2019); Reinhard and Nazarewicz (2021); Scamps et al. (2021).
@@
Except for doubly-magic nuclei which exhibit spherical shape, most atomic nuclei are deformed. Based on the RDFT and treating the effects of deformation, pairing correlations and continuum simultaneously, the deformed relativistic Hartree-Bogoliubov theory in continuum (DRHBc) was developed Zhou et al. (2010); Li et al. (2012a). Inheriting the advantages of the RCHB theory and including the deformation degrees of freedom, the DRHBc theory has been successfully applied in investigating the halo structures in B Yang et al. (2021); Sun (2021), C Sun et al. (2018, 2020); Wang et al. (2024), Ne Zhong et al. (2022); Pan et al. (2024), Na Zhang et al. (2023a), Mg Zhang et al. (2023b); An et al. (2024b), and Al Zhang et al. (2024a) isotopes, dripline locations In et al. (2021); Zhang et al. (2021); Pan et al. (2021); He et al. (2021, 2024), nuclear shapes and sizes Pan et al. (2019); Choi et al. (2022); Kim et al. (2022); Guo et al. (2023); Mun et al. (2023), nuclear stability and radioactive decays Xiao et al. (2023); Choi et al. (2024); Zhang et al. (2024b), as well as other applications and extensions Zhang et al. (2019); Sun and Zhou (2021a, b); Sun et al. (2022); Sun and Meng (2022); Zhang et al. (2022b, 2023c); Zhao et al. (2023); Mun et al. (2024); Zhang et al. (2024c); Pan et al. (2025). Based on the DRHBc theory and one of the most successful relativistic density functionals PC-PK1 Zhao et al. (2010), a nuclear mass table including both deformation and continuum effects is under construction Zhang et al. (2020); Pan et al. (2022). Up to now, the even-$Z$ part of the DRHBc mass table has been completed, and the ground-state properties, including the binding energy, charge radius and quadrupole deformation, have been published Zhang et al. (2022a); Guo et al. (2024a).
In this work, the systematics of the charge radii for even-$Z$ nuclei based on the DRHBc theory with the relativistic density functional PC-PK1 Zhao et al. (2010) are investigated. In Sec. II, a brief framework of the DRHBc theory is introduced. The numerical details in the DRHBc mass table calculations are provided in Sec. III. The systematics of the charge radii and their deformation correlation are discussed in Sec. IV. A summary is given in Sec. V.
@@ II Theoretical framework
where $\braket{r^{2}}_{p}$ is the root-mean-square (rms) radius of proton. Since Eq. (3) has been adopted in the fitting of the functional, the neutron contribution and the Darwin-Foldy correction Ekström et al. (2015) are not taken into account here.
@@ IV.1 Systematics of the charge radii
Figure 1(b) shows the charge radius differences between the DRHBc Guo et al. (2024a) and RCHB Xia et al. (2018) calculations, $R_{\text{ch}}^{\text{(DRHBc)}}-R_{\text{ch}}^{\text{(RCHB)}}$, for even-$Z$ nuclei. The rms difference is $\sigma=0.055$ fm. For nuclei near the regions with either proton or neutron magic numbers, the differences are negligible. Remarkable differences appear in the middle of the shell, for example, the regions with $N$ between 50 and 82, 82 and 126, 126 and 184, as well as 184 and 258. Among 4609 nuclei, there are 2808 nuclei with $R_{\text{ch}}^{\text{(DRHBc)}}-R_{\text{ch}}^{\text{(RCHB)}}\geqslant 0.01$ fm and only 67 nuclei with $R_{\text{ch}}^{\text{(DRHBc)}}-R_{\text{ch}}^{\text{(RCHB)}}\leqslant-0.01$ fm. This indicates that the deformation typically leads to a larger radius. Such effects are closely related to the shell structures. The details of the underlying mechanisms are discussed in Section IV.3.
@@
Figure 2: The ratio of $R_{\mathrm{ch}}$ calculated in DRHBc to that in RCHB as a function of $\beta_{2}$. The neutron numbers are scaled by colors. The empirical formula $R_{\mathrm{ch}}(\beta_{2})/R_{\mathrm{ch}}(0)=(1+5\beta_{2}^{2}/4\pi)^{1/2}$ is shown by the blue dashed line.
It would be interesting to systematically compare the charge radii between the prolate and oblate cases. In Fig. 2, the ratio of $R_{\mathrm{ch}}$ calculated in DRHBc to that in RCHB as a function of $\beta_{2}$ is shown. The blue dashed line corresponds to the empirical formula $R_{\text{ch}}(\beta_{2})=\left(1+\frac{5}{4\pi}\beta_{2}^{2}\right)^{1/2}R_{\text{ch}}(0)$. At $\beta_{2}\approx 0$, the ratios in the prolate and oblate sides exhibit similar behaviors, both close to the empirical curve. With the increase of $|\beta_{2}|$, the number of prolate nuclei is significantly larger than that of oblate nuclei. One cannot simply conclude that the radius is larger in the prolate or oblate side, as it varies significantly from case to case.
@@ IV.2 Evolution of the charge radius with the neutron number
To investigate the evolution of the charge radius with respect to the neutron number and the roles played by the deformation in detail, the Mo ($Z=42$) isotopes, which exhibit rich information on shape evolution as shown in Fig. 1(c), are taken as examples. The charge radii of Mo isotopes calculated by the DRHBc theory from the proton drip line at $N=40$ to the neutron drip line at $N=112$ are shown in Fig. 3(a) with red diamonds. For the isotopes with $54\leqslant N\leqslant 68$, two minima at the prolate and oblate sides are obtained with the energy differences ranging down to 0.1 MeV. Such distinct shapes with similar energies in the same nucleus suggest possible shape coexistence. Therefore, both the prolate and oblate solutions in the isotopes with $54\leqslant N\leqslant 68$ are presented with green and orange triangles, respectively. For comparison, the RCHB results Xia et al. (2018) are shown with blue circles, the data available Angeli and Marinova (2013); Li et al. (2021) are shown with black squares, and the empirical formula $r_{0}A^{1/3}$ is shown with the dashed line, where $r_{0}=0.956$ fm is determined by the data at $N=50$. In Fig. 3(a), the empirical formula roughly reproduces the general trend of the data. For the isotopes near $N=50$ with spherical or near-spherical shapes, both the RCHB and DRHBc calculations reproduce the data well, including the kink at $N=50$ corresponding to the shell closure. For the isotopes with $N>52$, the data are underestimated by RCHB but well reproduced by DRHBc. For the isotopes with shape coexistence, $58\leqslant N\leqslant 66$, the charge radii corresponding to the prolate minima in the DRHBc calculations better reproduce the data. The rms deviation from the data is $\sigma=0.030$ fm for DRHBc, in comparison with $\sigma=0.062$ fm for RCHB. In short, taking into account the deformation effect remarkably improves the description of charge radii for Mo isotopes.
@@ IV.3 Deformation dependence of the charge radius
Figure 5: (Color online) Total energies $E_{\text{tot}}$ [(a), (b) and (c)] and charge radii $R_{\text{ch}}$ [(d), (e) and (f)] as functions of quadrupole deformation $\beta_{2}$ of the Mo isotopes with $N=52,58,68$ from the constrained DRHBc calculations with PC-PK1. The ground-state deformations are denoted by squares. The empirical formula for charge radius, $R_{\text{ch}}(\beta_{2})=(1+5\beta_{2}^{2}/4\pi)^{1/2}R_{\text{ch}}(0)$, is shown by the gray dashed line.
To further investigate the deformation dependence of the charge radius, the shape coexistence and the competition between the prolate and oblate minima, taking the Mo isotopes with $N=52,58$ and 68 as examples, the potential energy curves (PECs), as well as the evolutions of the charge radii with the quadrupole deformation are shown in Fig. 5. The empirical formula for the charge radius, $R_{\text{ch}}(\beta_{2})=\left(1+\frac{5}{4\pi}\beta_{2}^{2}\right)^{1/2}R_{\text{ch}}(0)$ is also shown in Fig. 5.
@@
For the isotope with $N=68$, the ground state corresponds to the oblate minimum with $\beta_{2}=-0.23$, and there is a spherical minimum with the excitation energy 0.70 MeV, as shown in Fig. 5(c). For the neighboring isotopes with $N=69$ and 70, the ground states will move to the near-spherical and spherical shapes, respectively. The charge radius increases monotonically with the absolute value of the quadrupole deformation parameter, as shown in Fig. 5(f). It seems that the charge radius always increases with the deformation. The dependence of the charge radius on the deformation might provide a way to characterize the nuclear shape by a radius measurement, as has been discussed for triaxial nuclei in Ref. Maass et al. (2025).
@@
It is interesting to check whether the charge radii may decrease with the deformation $\beta_{2}$. In Fig. 6, the PEC, as well as the evolution of the charge radius with the quadrupole deformation are shown for 346Sg. In Fig. 6(a), there are two minima of the energy with $\beta_{2}=-0.15$ and $0.17$, and the ground state corresponds to the prolate one. The spherical result is a saddle point on the PEC, and its energy is 6.00 MeV higher than the ground state. In Fig.  6(b), there are two minima of the charge radius with $\beta_{2}=-0.15$ and $0.10$, which do not correspond to the minima in Fig. 6(a). Between these two minima of the charge radius, the charge radius increases with the decrease of the absolute value of the quadrupole deformation parameter. These results demonstrate that one cannot simply conclude the charge radius as a monotonically increasing function of the quadrupole deformation parameter, and the largest binding energy does not necessarily correspond to the lowest charge radius.
Figure 7: (Color online) The occupation probability $v_{i}^{2}$ and rms radius $r_{i}$ versus the single-proton energy $\epsilon_{i}$ near the Fermi energy $\lambda_{p}$ in the canonical basis for the deformed ground states of 100Mo [(a) and (b)] and 346Sg [(c) and (d)] with red solid lines. The corresponding spherical results are shown with blue dotted lines for comparison. In the spherical case, the levels are labeled by the spherical quantum numbers. In the deformed case, the levels stemming from $1g_{9/2}$ in 100Mo and $1i_{13/2}$ in 346Sg are labeled by the quantum numbers $m^{\pi}$. The proton Fermi energies $\lambda_{p}$ of the ground states are shown with green dashed lines.
To understand the underlying structural cause of the deformation effects on the charge radius, the composition of the rms radii for proton will be discussed. In Fig. 7, taking the deformed ground states of 100Mo and 346Sg as examples, the occupation probability $v_{i}^{2}$ and rms radius $r_{i}$ for the single-proton levels near the Fermi energy $\lambda_{p}$ in the canonical basis are shown with red solid lines. The corresponding spherical results are shown with blue dotted lines for comparison.
For 100Mo shown in Fig. 7(a) and (b), the inclusion of deformation leads to the splitting of single-proton levels, which significantly impacts the $v_{i}^{2}$ near $\lambda_{p}$ in Fig. 7(a), but the influence on the corresponding rms radius is small in Fig. 7(b). For the spherical case, the level $1g_{9/2}$ above the Fermi energy has the largest $r_{i}$, but small $v_{i}^{2}$. For the deformed case, this level $1g_{9/2}$ splits into $9/2^{+}$ and $7/2^{+}$ below the Fermi energy and $5/2^{+}$, $3/2^{+}$ and $1/2^{+}$ above the Fermi energy. The levels $9/2^{+}$ and $7/2^{+}$ have similar $r_{i}$ but large $v_{i}^{2}$, resulting in the larger $R_{\mathrm{ch}}$ for the deformed case in 100Mo.
For 346Sg shown in Fig. 7(c) and (d), the inclusion of deformation leads to the splitting of single-proton levels. For the spherical case, the level $1i_{13/2}$ below the Fermi energy has the largest $r_{i}$ and large $v_{i}^{2}$. For the deformed case, this level $1i_{13/2}$ splits into $1/2^{+}$, $3/2^{+}$, $5/2^{+}$, $7/2^{+}$ and $9/2^{+}$ below the Fermi energy and $11/2^{+}$ and $13/2^{+}$ above the Fermi energy. The levels $11/2^{+}$ and $13/2^{+}$ have similar $r_{i}$ but small $v_{i}^{2}$, resulting in the smaller $R_{\mathrm{ch}}$ for the deformed case in 346Sg.
The deformation dependences of the charge radii in 100Mo and 346Sg show that the increase or decrease of the charge radii with deformation is related to specific shell structures, particularly the key single-particle levels near the Fermi energy.
@@ V Summary
In the present, only the results of even-$Z$ nuclei are available Guo et al. (2024a). A complete DRHBc mass table including odd-$Z$ nuclei will appear soon, enabling a more comprehensive investigation on the systematics of the charge radii and their deformation correlation. More experimental data, including charge radii and deformations, will deepen our understanding on nuclear structures, especially the data distinguishing prolate from oblate shapes, which will provide valuable constraints in examining existing nuclear models. It would also be interesting to use DRHBc to improve the isospin-dependent formula for the charge radius in Ref. Zhang et al. (2002b) by incorporating the deformation effect. As the DRHBc theory is based on the axial symmetry approximation, the effects of the triaxial deformation Zhang et al. (2023d) and the deformations with higher orders on the charge radius are issues worthy of further investigation.